The demand and supply equations for a product are: Qd = 300 – 6P and Qs = -40 + 6P. • Determine the market equilibrium and draw graphs. • Suppose that the government decides to impose a flat tax of 10% on each unit sold. Show that the price that consumers pay would be the same if the government imposed a tax of Rs. 1.70 per unit sold. Draw graphs and explain. • Also calculate the total revenue earned by sellers before and after the tax, the tax revenue raised by the government, changes in consumer and producers surplus and dead weight loss. 


1
Expert's answer
2022-02-21T11:56:30-0500

At market equilibrium, "Qd=Qs"

"300-6P=-40+6P"

"12P=340"

"P=28.3"

"Q=300-6\u00d728.3=130.2"

Taxes

10% of 28.3 is 2.83 Rs compared to a tax of 1.70 Rs per unit, the reason why the consumers will pay same prices for the tax scenarios is because in the first case, the extra tax burden will be distributed evenly between the producers and the consumers while in the second case the whole tax burden will be shouldered by the consumer.

Revenues

Total Revenue before tax:

"130.2\u00d728.3=3685Rs"

Total Revenue after tax:

"TR=130.2\u00d7(28.3+2.83)=4053"

Total tax Revenue for the government:

"TR_t=130.2\u00d72.83=368.5"